OpenStudy (anonymous):

A type of rope weighs 4 ounces per inch. What does it weigh in pounds per foot?

OpenStudy (tkhunny):

\(\dfrac{4\;oz}{1\;in}\cdot\dfrac{1\;lb}{16\;oz}\cdot\dfrac{12\;in}{1\;ft}\) Please examine each fraction carefully and see what it does.

OpenStudy (ja1):

@hdmartin7 in Laymen terms: You need to get 1 inch to 12 since 12 inches = 1 foot and 4 oz to 16 oz since it is 1 lb